OpenStudy (anonymous):

Please help, I'm so confused by this problem! (cos x - sqrt 2/2)(sec x -1)=0 Part1: Use the zero product property to set up two equations that will lead to solutions to the original equation. Part 2: Use a reciprocal identity to express the equation involving secant in terms of sine, cosine, or tangent. Part 3: Solve each of the two equations in Part 1 for x, giving all solutions to the equation.

OpenStudy (zzr0ck3r):

a*b = 0 either a = 0, b = 0 , or both a and b are = 0 so find out what will make the last happen to cover all bases cos(x) - sqrt(2)/2 = 0 so cos(x) = sqrt(2)/2 and (sec x -1)=0 so sec(x) = 1 1/cos(x) = 1 cos(x) = 1

OpenStudy (zzr0ck3r):

you need to solve those two equations, can you do that?

OpenStudy (zzr0ck3r):

there is going to be infinite solutions

OpenStudy (anonymous):

Thank you so much! :D